# Dual Prompt Tuning for Domain-Aware Federated Learning

### v1

@@ Abstract
Federated learning is a distributed machine learning paradigm that allows multiple clients to collaboratively train a shared model with their local data. Nonetheless, conventional federated learning algorithms often struggle to generalize well due to the ubiquitous domain shift across clients. In this work, we consider a challenging yet realistic federated learning scenario where the training data of each client originates from different domains. We address the challenges of domain shift by leveraging the technique of prompt learning, and propose a novel method called Federated Dual Prompt Tuning (Fed-DPT). Specifically, Fed-DPT employs a pre-trained vision-language model and then applies both visual and textual prompt tuning to facilitate domain adaptation over decentralized data. Extensive experiments of Fed-DPT demonstrate its significant effectiveness in domain-aware federated learning. With a pre-trained CLIP model (ViT-Base as image encoder), the proposed Fed-DPT attains 68.4% average accuracy over six domains in the DomainNet dataset, which improves the original CLIP by a large margin of 14.8%.
1 Introduction
Federated Learning (FL) is a privacy-preserving machine learning technique that allows training a centralized model across decentralized devices while keeping data localized. The general federated learning paradigm involves many rounds of local training and global parameter aggregation(McMahan et al., 2017), which enables learning from decentralized data but is susceptible to two primary challenges: extensive domain shift across clients and limited communication efficiency.
As federated learning assumes that each local dataset is independently collected, the data is inevitably heterogeneous across clients. This heterogeneity highly challenges FL algorithms and results in a considerable performance gap realative to centralized training. In practice, as it is difficult to collect a real federated dataset from multiple clients, prior works choose to partition one single dataset into several splits with non-independently and identically distributed (non-i.i.d.) labels so that to simulate the desired heterogeneity. However, we argue that this setting overlooks a crucial characteristic of federated datasets, which is that data from various clients may originate from different domains, and the data should exhibit significant differences in the input space, rather than simple non-i.i.d. characteristics based on labels.
@@ 1 Introduction
Therefore, we consider a more challenging yet realistic scenario: the clients desire to deal with the same machine learning problem (e.g., image classification with the same target categories), yet their local data originate from different domains. Following the practice of prior works (Peng et al., 2020), we formulate this scenario using domain-aware datasets like DomainNet (Peng et al., 2019), where there are labeled images sourced from six distinct domains with quite different styles such as real-world, paining, and sketch. Due to the large diversity in input, conventional domain-agnostic federated learning approaches often struggle to generalize well in this problem.
The extensive domain shift and data heterogeneity also challenge the convergence abilities of decentralized training, leading to non-robust federated learning models. This effect becomes more pronounced when dealing with larger models. For example, it is observed that in some federated learning scenarios with highly heterogeneous data, a 152-layer ResNet (He et al., 2016) with 60M parameters even performs worse than a 50-layer ResNet with 26M parameters (Qu et al., 2022). A potential way to tackle this challenge is leveraging parameter-efficient training strategies, which employ a pre-trained model and only fine-tune a small portion of parameters (Lu et al., 2023).
With the advancement of Contrastive Language-Image Pre-training (CLIP) models (Radford et al., 2021), it is very convenient to develop parameter-efficient learning protocols by prompt tuning techniques. Specifically, these methods employ a pre-trained CLIP model, freezing both its image and text encoders, and feed it with learnable tokens (i.e., prompts) attached to the original input (Zhou et al., 2021; Jia et al., 2022). By optimizing only the prompt tokens, the model can quickly adapt to downstream datasets and domains.
In this work, we propose Federated Dual Prompt Tuning (Fed-DPT), a novel federated learning approach to overcome the challenges mentioned above. In detail, we address the challenge of parameter-efficiency by harnessing the techniques of CLIP and prompt learning for both visual and textual inputs, making our method friendly to communication cost and robust to federated optimization with heterogeneous data. While existing CLIP-based methods share one single prompt learner (Guo et al., 2023) or adaptor (Lu et al., 2023) with all images from various domains, we further tackle the challenge of domain shift across clients by introducing domain-specific prompts and coupling visual and textual representations by self-attention, leading to a domain-aware federated learning method.
Extensive experiments demonstrate the significant effectiveness of our method. Remarkably, we obtain a 68.4% average accuracy over six domains in the DomainNet dataset, outperforming the original CLIP model by 14.8%. Compared with conventional federated learning methods such as FedAvg (McMahan et al., 2017) and FedProx Li et al. (2020b), and existing domain-agnostic CLIP-based approaches such as PromptFL (Guo et al., 2023) and FedCLIP (Lu et al., 2023), our Fed-DPT consistently achieves superior performance on three benchmarks.
2 Related Work
@@
Federated learning. The concept of federated learning was first introduced in the Federated Averaging (FedAvg) paper (McMahan et al., 2017). to address machine learning problems with massively distributed private data. To enhance the learning potential of FedAvg, FedProx (Li et al., 2020b) introduces a $\ell_{2}$ regularization term into the original federated learning objective. In addition, based on the success of FedAvg, many follow-up works improve federated learning in terms of privacy-preserving potentials (Wei et al., 2020; Truex et al., 2019), robustness to heterogeneous data (Karimireddy et al., 2020; Li et al., 2019), communication efficiency (Konečnỳ et al., 2016; Sattler et al., 2019), and compatibility to model architectures (Li et al., 2020a; Qu et al., 2022). In contrast to general federated learning methods that simulate non-i.i.d. data by partitioning datasets in the label space, many recent works consider federated learning in a more realistic context of domain adaptation (Yao et al., 2022; Shenaj et al., 2023; Peng et al., 2020). Recently, based on the advances in multi-modal contrastive learning (Radford et al., 2021), various works develop CLIP-based federated learning methods. For example, FedCLIP (Lu et al., 2023) uses a pre-trained CLIP model and performs federated training on an additional adaptor layer, and PromptFL (Guo et al., 2023) proposes to use prompt learning methods for federated optimization.
@@
Vision-language models. Following the success of contrastive pre-training in visual modality (He et al., 2020; Chen et al., 2020; Grill et al., 2020; Caron et al., 2021; Chen & He, 2021; Chen et al., 2021), multi-modal contrastive pre-training has become a common paradigm in recent years as well. A representative work is CLIP (Radford et al., 2021), which jointly pre-trains a visual and a textual encoder using an InfoNCE objective (Gutmann & Hyvärinen, 2010) with around 400 million curated image-text pairs. ALIGN (Jia et al., 2021) improves CLIP by scaling up the training dataset to 1.8 billion noisy image-text pairs, and BASIC (Pham et al., 2021) further increases the scale of both data and model. As a result, such CLIP-like models allow zero-shot inference when it comes to transfer learning on downstream tasks.
@@
Prompt tuning. While fine-tuning a pre-trained model for downstream machine learning tasks has traditionally dominated the field of transfer learning, recent progress in prompt learning offers a compelling alternative. Specifically, the prompt tuning techniques fine-tune learnable prompt tokens attached to CLIP’s inputs instead of training the entire model (Zhou et al., 2021; 2022; Wang et al., 2023; Yao et al., 2023). There also exist prompt tuning protocols for visual modality (Jia et al., 2022) and both visual and textual modalities (Yao et al., 2021; Zang et al., 2022). Similarly, there are adapter-based methods designed for CLIP-like models, which also freeze the encoders and only fine-tune several newly attached layers on top of them (Gao et al., 2021; Zhang et al., 2021).
3 Preliminaries
3.1 Contrastive Language-Image Models
@@
Contrastive Language-Image Pre-training (Radford et al., 2021) is a weakly supervised learning paradigm that combines visual and language encoders to solve image recognition problems. Formally, CLIP has an image encoder ${\bm{F}}_{V}:{\mathbb{R}}^{3\times w\times h}\rightarrow{\mathbb{R}}^{d}$ where $w$ and $h$ denotes the input image’s spatial resolution and $d$ denotes the dimension of the latent space, and a text encoder ${\bm{F}}_{T}:{\mathbb{R}}^{l\times d_{e}}\rightarrow{\mathbb{R}}^{d}$ where $l$ is the length of input sentence and $d_{e}$ is the dimension of word embedding (512 for CLIP’s transformer). CLIP is trained by image-text pairs, in which the text briefly describes the information in the image. By encoding both the image and text into the same latent space, CLIP can learn an alignment between visual and textual input with a contrastive loss (Gutmann & Hyvärinen, 2010).
The CLIP-like vision-language models are generally pre-trained on web-scale datasets with several hundred million image-text pairs (Radford et al., 2021; Jia et al., 2021; Pham et al., 2021). As a result, these models excel in recognizing image features and support zero-shot inference by aligning visual features to text queries. Specifically, given an image, the encoder ${\bm{F}}_{V}$ maps it into a vectorized feature ${\bm{f}}_{V}\in{\mathbb{R}}^{d}$, and given several text queries (i.e., class names) such as “cat”, “dog”, and “horse”, the encoder ${\bm{F}}_{T}$ maps each of them into a vector ${\bm{f}}_{T}\in{\mathbb{R}}^{d}$ with a prompt. By computing the cosine similarity between visual and textual features, CLIP classifies the image into the $i$-th class with probability
$p_{i}=\frac{\exp(<{\bm{f}}_{V},{\bm{f}}_{T}^{i}>/\tau)}{\sum_{j}\exp(<{\bm{f}}_{V},{\bm{f}}_{T}^{j}>/\tau)},$ (1)
where <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes dot product, $\tau$ denotes a temperature coefficient, and note that both ${\bm{f}}_{V}$ and ${\bm{f}}_{T}$ have been $\ell_{2}$ normalized.
3.2 Prompt Tuning for Vision and Language
@@
Despite CLIP’s impressive zero-shot inference capabilities, it still exhibits a noticeable accuracy gap in comparison to in-domain fine-tuning. However, tuning CLIP’s model parameters can easily break the well-established alignment between vision and language, and CLIP therefore loses the ability of open-vocabulary inference as well as address the challenge of domain adaptation. Instead, prompt tuning attaches learnable tokens to the input, leaving the feature encoders fixed. This approach allows the model to retain its zero-shot and open-set inference abilities while significantly improving in-domain accuracy.
@@
Textual Prompt Tuning (TPT). As previously mentioned, CLIP’s text query consists of a hand-crafted prompt (also referred to as prefix) such as “A photo of a” and a class name such as “dog”. TPT replaces the prefix by learnable vectors (Zhou et al., 2021). Formally, it feeds the text encoder ${\bm{F}}_{T}$ with a sequence of trainable vectors ${\bm{t}}_{1},\ {\bm{t}}_{2},\ \dots,\ {\bm{t}}_{m}$ followed by the embedding of class name, where each ${\bm{t}}\in{\mathbb{R}}^{d_{T}}$ represents a “latent word” that has the same dimension as CLIP’s word embedding. During training, both CLIP’s vision and language encoders are frozen and only the prompt vectors ${\bm{t}}_{1},\ {\bm{t}}_{2},\ \dots\ {\bm{t}}_{m}$ are optimized.
@@
Visual Prompt Tuning (VPT). The prompt tuning protocol also works for visual input if the image encoder is a transformer-like model such as the Vision Transformer (Dosovitskiy et al., 2021). Specifically, this method attaches $n$ trainable vectors ${\bm{i}}_{1},\ {\bm{i}}_{2},\ \dots,\ {\bm{i}}_{n}$ to the patch-wise embedded image, and uses an additional head to project the output, where each prompt token ${\bm{i}}\in{\mathbb{R}}^{d_{I}}$ has the same dimension as image embeddings (e.g., 768 for ViT-Base). In VPT, only the prompt tokens and the head are optimized.
4 Methodology
We present Federated Dual Prompt Tuning (Fed-DPT), a novel federated learning method to overcome the challenges of domain shift and parameter efficiency. We detail our method below.
@@ 4 Methodology
Figure 2: Parameter aggregation pipeline of Fed-DPT. We aggregate textual prompts by concatenating the domain-specific tokens from each client, and aggregate visual prompts by averaging.
4.1 Problem Formulation
There are $n$ clients that desire to deal with the same machine learning problem, e.g., image classification with the same target categories. The $n$ clients possess their own training data that originate from $n$ distinct domains. In other words, each client stands for a specific domain. We simulate this scenario using domain adaptation datasets like DomainNet (Peng et al., 2019), which encompass images from six different domains including clipart, information graph, painting, quickdraw, real-world images, and sketch. As the image features exhibit significant variation across different domains, it is indeed a challenging task for federated optimization. However, it is a realistic scenario because many times, the data heterogeneity between clients arises from differences in feature distributions rather than label distributions. Notably, our setting is compatible with the task that clients have non-i.i.d. labels. In our ablation study, we also further divide each domain into five splits with non-i.i.d. categories (see Section 5.3 for details).
4.2 Local training
With CLIP, a very simple way to deal with domain shift is to use domain-aware prompt contexts for text queries. For example, in DomianNet, when we use prefix “a painting of a” for the painting domain, and use “a sketch of a” for the sketch domain, the predictions can be more accurate and robust. This idea is also referred to as domain-specific prompts (Ge et al., 2022), while employing learnable text prompts can further improve the predictive performance. Inspired by this observation, we propose to use domain-specific prompts for CLIP’s text encoder. Formally, we define a text prompt by a sequence of learnable tokens:
@@ 4.2 Local training
Figure 1 illustrates our Fed-DPT’s local training framework. We initialize Fed-DPT by loading the same CLIP model for each client and freezing the parameters of both the image encoder ${\bm{F}}_{V}$ and the text encoder ${\bm{F}}_{T}$. For our task, we have $n$ text prompts ${\bm{P}}_{T}^{1},{\bm{P}}_{T}^{2},\dots,{\bm{P}}_{T}^{n}$ corresponding to the $n$ domains. During local training, the $n$ text prompts are shared among the clients, yet the $i$-th prompt ${\bm{P}}_{T}^{i}$ can only be trained by the $i$-th client (we will detail this mechanism later). We separately feed the encoder ${\bm{F}}_{T}$ with all the $n$ text prompts followed by a class name, leading to $n$ representation vectors ${\bm{f}}_{T}^{1},{\bm{f}}_{T}^{2},\dots,{\bm{f}}_{T}^{n}$, where
@@
Beside of ${\bm{f}}_{V}$, we meanwhile obtain attention scores between the [cls] token and visual prompts. Formally, denoting ${\bm{q}}_{\mbox{cls}}$ as the query vector of the class token, and ${\bm{k}}_{i}$ as the key vector of the $i$-th prompt token in ${\bm{F}}_{V}$’s last self-attention block, we have ${\bm{w}}=[w_{1},w_{2},\dots,w_{n}]$ with
@@
During the local training process in the $i$-th client, we optimize the $i$-th text prompt ${\bm{P}}_{T}^{i}$ and all the $n$ visual prompts $[{\bm{v}}]_{1},[{\bm{v}}]_{2},\dots,[{\bm{v}}]_{n}$ by a $\ell_{2}$ loss applied to ${\bm{f}}_{V}$ and ${\bm{f}}_{T}$:
${\mathcal{L}}=<{\bm{f}}_{V},{\bm{f}}_{T}>/||{\bm{f}}_{V}||\cdot||{\bm{f}}_{T}||.$ (7)
Here we explain why we optimize these parameters. We desire the $i$-th text prompt ${\bm{P}}_{T}^{i}$ to represent the features of the $i$-th domain in the latent space of textual embeddings. However, the $i$-th client only possesses images from the $i$-th domain, so we cannot train ${\bm{P}}_{T}^{j}\ (j\neq i)$ yet instead load them from other clients. We introduce visual prompts to detect the correlations between an input image and the $n$ domains, so it is fine to optimize all of them. A detailed comparison of different training strategies can be found in our ablation study (see Table 4a and 4b for details). Also note that in our experiments, we find the $\ell_{2}$ loss yields better predictive performance and allows more flexible training compared with cross-entropy, while we do not observe any collapse issues.
4.3 Parameters Aggregation
As mentioned above, for the $i$-th client, we optimize ${\bm{P}}_{T}^{i}$ by gradients and load ${\bm{P}}_{T}^{j}\ (j\neq i)$ from other clients, so the aggregation of text prompts does not involve parameter merging processes (e.g. averaging). We illustrate the aggregation pipeline in Figure 2. For easy understanding, we suppose there is a centralized parameter server — actually, Fed-DPT also works for decentralized communication — and the clients upload their corresponded text prompt to it in each communication round. The sever concatenates the $n$ uploaded text prompts and then sends to every client. For visual parameters, as all visual prompts are optimized by every client, we perform federated averaging in the server and then send the merged parameters to each client. Note that we do not need to share CLIP encoders’ parameters as each client is initialized with the same CLIP model and its parameters are frozen during training.
This parameter aggregation paradigm works well for Fed-DPT, yet may create a minor problem for the text encoder. Specifically, after each communication round, the external text prompts of the $i$-th client, i.e., ${\bm{P}}_{T}^{j}\ (j\neq i)$ will be re-loaded. We observe that this sudden change of parameters often negatively affects our model. To address this issue, we propose to apply momentum update (also referred to as exponential moving average) to the external text prompts. Formally, we have
@@ 4.3 Parameters Aggregation
5 Experiments
5.1 Experimental Setup
Datasets. We evaluate our Fed-DPT and baseline methods on the following three domain adaptation image classification benchmarks:
  • DomainNet (Peng et al., 2019). The DomainNet dataset has around 600,000 images spanning 345 categories from six domains, which covers diverse image styles including clipart, infograph, painting, quickdraw, real, and sketch.
  • OfficeHome (Venkateswara et al., 2017). The OfficeHome dataset consists of approximately 15,500 images depicting everyday objects in 65 classes. It further categorizes the images into four domains: art, clipart, product, and real-world.
  • PACS (Li et al., 2017). The PACS dataset contains around 10,000 images drawn from seven categories and four domains, including photo, sketch, cartoon, and painting styles.

### v5

# Learning to Prompt Your Domain for Vision-Language Models

@@ Abstract
Prompt learning has recently become a very efficient transfer learning paradigm for Contrastive Language Image Pretraining (CLIP) models. Compared with fine-tuning the entire encoder, prompt learning can obtain highly competitive results by optimizing only a small number of parameters, which presents considerably exciting benefits for federated learning applications that prioritizes communication efficiency. However, in this work, we identify that directly transferring prompt learning approaches into federated learning does not yield favorable results since the model often suffers from considerable domain gaps across different clients. To address this issue, we propose ADAPT, a novel domain-aware prompt learning approach that facilitates both intra- and inter-domain prompts across federated participants. The basic idea of ADAPT is that the prompted CLIP should detect the input image’s domain correspondence and before making the prediction of its category. Extensive experiments of ADAPT demonstrate its significant efficiency and effectiveness in federated learning. For example, by learning and sharing only 0.08M parameters, our ADAPT attains a 68.4% average accuracy over six domains in the DomainNet dataset, which improves the original CLIP by a large margin of 14.8%.
Figure 1: An intuitive comparison of our ADAPT to zero-shot CLIP and the basic domain-agnostic prompt learning approach.
Introduction
Contrastive Language Image Pretraining (CLIP) (Radford et al. 2021) has recently been proven to be a powerful model for multi-modal representation learning. By connecting the feature spaces of images and texts, it enables convenient open-vocabulary classification simply by matching the visual representations of images with the textual embeddings of their class names. Building upon CLIP, the prompt learning technique, which freezes all encoders while introducing learnable tokens at the input side of the model, can help the model adapt to downstream domains with minimal cost.
Compared to the traditional finetuning paradigm, prompt learning techniques built on CLIP can achieve highly competitive results with a minimal amount of learnable parameters (e.g., 0.1% of encoder parameters). This significant advantage in parameter efficiency motivates us to explore the application of prompt-based CLIP in federated learning — In federated learning, participants need to frequently share and update the model’s learnable parameters, leading to high communication costs and slow convergence rates, while prompt learning offers a highly parameter-efficient approach to fundamentally address these issues.
@@ Introduction
In this work, we consider a challenging yet realistic federated learning scenario: the participants aim to deal with the same machine learning problem (e.g., image classification with the same target categories), yet their local data originate from different domains. Following prior practice (Peng et al. 2020), we formulate this scenario using domain-aware datasets like DomainNet (Peng et al. 2019), where there are labeled images sourced from six distinct domains with quite different styles such as real-world, paining, and sketch. Due to the large diversity in input, conventional domain-agnostic federated learning approaches often struggle to generalize well in this problem.
It is noteworthy that this federated learning scenario is more aligned with real-world conditions, as the data heterogeneity among different participants often manifests as variations in feature distributions and image styles, rather than the imbalanced label distributions that most empirical studies use for simulating a non-independent and identical distribution (non-IID). However, in this scenario, due to significant domain gaps between participants, directly applying conventional prompt learning methods cannot yield satisfactory results. We present an intuitive comparison between our method and the existing approaches in Figure 1
To address this issue, we propose a simple but highly effective approach called FederAted Domain-Aware Prompt Tuning (ADAPT). In detail, our ADAPT approach specifically set up a visual and a textual prompt for each potential domain. Each text prompt consists of several learnable tokens that represent textual descriptions indicative of each domain’s style information. Each visual prompt denotes a single learnable token that is appended to patch-embedded image tokens. We optimize the prompts through two loss functions: 1) a basic object classification loss, which is applied between the global feature of the input image and the textual representation of its corresponding class name; 2) a domain correspondence loss, which is applied between each pair of visual and textual prompts’ output. A detailed framework of our method is illustrated in Figure 2
The inference process of ADAPT involves two steps. First, given an input image, the vision encoder can determine the probability or confidence level that the image belongs to each given domain by measuring the attention scores to each visual prompt token. Next, based on these domain correspondence probabilities, we perform cross attention on the representation vectors of the given class name under each textual prompt to produce its final text feature. By finding the text feature that best matches the image feature, we can determine the classification result of the image.
Following extensive theoretical analyses and empirical evaluations, we have identified several significant advantages of ADAPT, which serve as our main contributions:
  • Domain-aware prompt learning. We assign domain-specific textual prompts to each federated participant, enabling the model to make predictions by input images’ corresponding domain information, which effectively addresses the widespread issue of domain gaps in federated learning. Our experimental results demonstrate a significant performance improvement with this design: based on a pretrained CLIP model equipped with a ViT-Base image encoder, our ADAPT method achieves an average accuracy of 68.4% on DomainNet, which significantly outperforms the zero-shot CLIP’s 53.6%, the basic prompt learning’s 63.2%, and PedProx’s (Li et al. 2020) 55.3% with a same image encoder.
  • Efficient communication. Our ADAPT approach requires training and sharing only a small fraction of parameters, significantly reducing the communication overhead in federated learning. For instance, in our domain-aware federated learning experiments, ADAPT, with just 0.08M trainable parameters, achieved state-of-the-art performance on the DomainNet (Peng et al. 2019), OfficeHome (Venkateswara et al. 2017), and PACS (Li et al. 2017) datasets.
  • Superior privacy preservation. Due to the minimal amount of trainable parameters in ADAPT, traditional federated learning attack algorithms (Zhu, Liu, and Han 2019; Geiping et al. 2020) struggle to reconstruct the local data of participants from model gradients. Additionally, the learnable prompts themselves do not leak customer privacy—we have attempted to decode our prompts but found it difficult to extract any interpretable information from them.
Related Work
@@
Federated learning was first introduced in the Federated Averaging (FedAvg) paper (McMahan et al. 2017), addressing machine learning problems with massively distributed private data. To enhance the learning potential of FedAvg, FedProx (Li et al. 2020) adds a $\ell_{2}$ regularization term into the original federated learning objective. Following FedAvg’s success, many follow-up works improve federated learning in terms of privacy-preserving potentials (Wei et al. 2020; Truex et al. 2019), robustness to heterogeneous data (Karimireddy et al. 2020; Li et al. 2019), communication efficiency (Konečnỳ et al. 2016; Sattler et al. 2019), and compatibility to model architectures (Li, Wen, and He 2020; Qu et al. 2022). In contrast to general federated learning methods that simulate non-i.i.d. data by partitioning datasets in the label space, many recent works consider federated learning in a more realistic context of domain adaptation (Yao et al. 2022; Shenaj et al. 2023; Peng et al. 2020). Recently, based on advances in multi-modal contrastive learning (Radford et al. 2021), various works develop CLIP-based federated learning methods. For example, FedCLIP (Lu et al. 2023) uses a pre-trained CLIP model and performs federated training on an additional adaptor layer, and PromptFL (Guo et al. 2023) proposes to use prompt learning methods for federated optimization.
@@
Vision-language models. Following the success of contrastive pre-training in visual modality (He et al. 2020; Chen et al. 2020; Grill et al. 2020; Caron et al. 2021; Chen and He 2021; Chen, Xie, and He 2021), multi-modal contrastive pre-training has become a common paradigm in recent years as well. A representative work is CLIP (Radford et al. 2021), which jointly pre-trains a visual and a textual encoder using an InfoNCE objective (Gutmann and Hyvärinen 2010) with around 400 million curated image-text pairs. ALIGN (Jia et al. 2021) improves CLIP by scaling up the training dataset to 1.8 billion noisy image-text pairs, and BASIC (Pham et al. 2021) further increases the scale of both data and model. As a result, such CLIP-like models allow zero-shot inference when it comes to transfer learning on downstream tasks.
@@
Prompt tuning. While fine-tuning a pre-trained model for downstream machine learning tasks has traditionally dominated the field of transfer learning, recent progress in prompt learning offers a compelling alternative. Specifically, the prompt tuning techniques fine-tune learnable prompt tokens attached to CLIP’s inputs instead of training the entire model (Zhou et al. 2021, 2022; Wang et al. 2023; Yao, Zhang, and Xu 2023). There also exist prompt tuning protocols for visual modality (Jia et al. 2022) and both visual and textual modalities (Yao et al. 2021; Zang et al. 2022). Similarly, there are adapter-based methods designed for CLIP-like models, which also freeze the encoders and only fine-tune several newly attached layers on top of them (Gao et al. 2021; Zhang et al. 2021).
Preliminaries
Contrastive Language-Image Pre-training (CLIP)
@@
CLIP is a weakly supervised learning paradigm that combines visual and language encoders to solve image recognition problems. Formally, CLIP has an image encoder ${\bm{F}}_{V}:{\mathbb{R}}^{3\times w\times h}\rightarrow{\mathbb{R}}^{d}$ where $w$ and $h$ denotes the input image’s spatial resolution and $d$ denotes the dimension of the latent space, and a text encoder ${\bm{F}}_{T}:{\mathbb{R}}^{l\times d_{e}}\rightarrow{\mathbb{R}}^{d}$ where $l$ is the length of input sentence and $d_{e}$ is the dimension of word embedding. CLIP is trained by image-text pairs, in which the text briefly describes the information in the image. By encoding both image and text into the same latent space, CLIP can learn an alignment between visual and textual input with a contrastive loss (Gutmann and Hyvärinen 2010). During inference, CLIP supports zero-shot classification by match the visual representation of input image and the textual representation of target class names.
Prompt Tuning for Vision and Language
@@
Despite CLIP’s impressive zero-shot inference capabilities, there remains a noticeable accuracy gap in comparison to in-domain fine-tuning. However, fine-tuning the CLIP model may easily break the well-established alignment between vision and language, and CLIP will therefore loses the ability of open-vocabulary inference. Instead, prompt tuning attaches learnable tokens to the input, leaving the feature encoders fixed, which allows the model to retain its zero-shot and open-set inference abilities while significantly improves in-domain accuracy.
@@
Textual Prompt Tuning (TPT). As previously mentioned, CLIP’s text query consists of a hand-crafted prompt (also referred to as prefix) such as “A photo of a” and a class name such as “dog”. TPT replaces the prefix by learnable vectors (Zhou et al. 2021). During training, both CLIP’s vision and language encoders are frozen and only the prompt vectors are optimized.
@@
Visual Prompt Tuning (VPT). The prompt tuning protocol also works for visual input if the image encoder is a transformer-like model such as the Vision Transformer (Dosovitskiy et al. 2021). Specifically, this method attaches trainable vectors to the patch-wise embedded image, and uses an additional head to project the output. In VPT, only the prompt tokens and the head are optimized.
@@
Methodology
Problem Formulation. Supposing there are $n$ clients that desire to deal with the same machine learning problem, e.g., image classification with the same target categories. The $n$ clients possess their own training data that originate from $n$ distinct domains. In other words, each client stands for a specific domain. We simulate this scenario using domain adaptation datasets like DomainNet (Peng et al. 2019), which encompass images from six different domains including clipart, information graph, painting, quickdraw, real-world images, and sketch. As the image features exhibit significant variation across different domains, it is indeed a challenging task for federated optimization. However, it is a realistic scenario because many times, the data heterogeneity between clients arises from differences in feature distributions rather than label distributions. Notably, our setting is compatible with the task that clients have non-i.i.d. labels. In our ablation study, we also further divide each domain into five splits with non-i.i.d. categories.
Local training
With CLIP, a very simple way to deal with domain shift is to use domain-aware prompt contexts for text queries. For example, in DomianNet, when we use prefix “a painting of a” for the painting domain, and use “a sketch of a” for the sketch domain, the predictions can be more accurate and robust. This idea is also referred to as domain-specific prompts (Ge et al. 2022), while employing learnable text prompts can further improve the predictive performance. Inspired by this observation, we propose to use domain-specific prompts for CLIP’s text encoder. Formally, we define a text prompt by a sequence of learnable tokens:
@@
Figure 2 illustrate our ADAPT’s local training framework and process. We initialize ADAPT by loading the same CLIP model for each client and freezing the parameters of both the image encoder ${\bm{F}}_{V}$ and the text encoder ${\bm{F}}_{T}$. For our task, we have $n$ text prompts ${\bm{P}}_{T}^{1},{\bm{P}}_{T}^{2},\dots,{\bm{P}}_{T}^{n}$ corresponding to the $n$ domains. During local training, the $n$ text prompts are shared among the clients, yet the $i$-th prompt ${\bm{P}}_{T}^{i}$ can only be trained by the $i$-th client (we will detail this mechanism later). We separately feed the encoder ${\bm{F}}_{T}$ with all the $n$ text prompts followed by a class name, leading to $n$ representation vectors ${\bm{f}}_{T}^{1},{\bm{f}}_{T}^{2},\dots,{\bm{f}}_{T}^{n}$, where
@@
We obtain the final textual representation through a cross attention layer. To minimize the number of parameters as much as possible, here we replace the query, key, and value projection matrices in the cross attention block with identity matrices, which we empirically find does not affect performance too much. Formally, denoting ${\bm{q}}_{\mbox{cls}}$ as the query vector of the class token, and ${\bm{k}}_{i}$ as the key vector of the $i$-th prompt token in ${\bm{F}}_{V}$’s last self-attention block, we have ${\bm{w}}=[w_{1},w_{2},\dots,w_{n}]$ with
@@
During training, we optimize the model (actually learnable parameters only appear in prompts) by an object classification loss, which is a cross-entropy function applied between ${\bm{f}}_{V}$ and ${\bm{f}}_{T}$, and a domain correspondence loss which is another cross entropy function applied between each pair of visual and textual outputs. Here we explain why we optimize these parameters. We desire the $i$-th text prompt ${\bm{P}}_{T}^{i}$ to represent the features of the $i$-th domain in the latent space of textual embeddings. However, the $i$-th client only possesses images from the $i$-th domain, so we cannot train ${\bm{P}}_{T}^{j}\ (j\neq i)$ yet instead load them from other clients. We introduce visual prompts to detect the correlations between an input image and the $n$ domains, so it is fine to optimize all of them. A detailed comparison of different training strategies can be found in our ablation study (see Table 5a and 5b for details).
Parameters Aggregation
As mentioned above, for the $i$-th client, we optimize ${\bm{P}}_{T}^{i}$ by gradients and load ${\bm{P}}_{T}^{j}\ (j\neq i)$ from other clients, so the aggregation of text prompts does not involve parameter merging processes (e.g. averaging).
Suppose there is a centralized parameter server — although ADAPT also works for decentralized communication — and the clients upload their corresponded text prompt to it in each communication round. The server concatenates the $n$ uploaded text prompts and sends to every client. For visual parameters, as all visual prompts are optimized by every client, we perform federated averaging in the server and then send the merged parameters to each client. Note that we do not need to share CLIP encoders’ parameters as each client is initialized with the same CLIP model and its parameters are frozen during training.
This parameter aggregation paradigm works well for ADAPT, yet may create a minor problem for the text encoder. Specifically, after each communication round, the external text prompts of the $i$-th client, i.e., ${\bm{P}}_{T}^{j}\ (j\neq i)$ will be re-loaded. We observe that this sudden change of parameters often negatively affects our model. To address this issue, we propose to apply momentum update (also referred to as exponential moving average) to the external text prompts. Formally, we have
@@
Privacy Preservation
In ADAPT, there are two potential ways to expose participants‘ private data. First, similar to most federated learning algorithms, our ADAPT requires to share gradient information across all participants, so some private information might be able to be reconstructed by gradient-based attacking algorithms such as Deep Leakage from Gradient (DLG) (Zhu, Liu, and Han 2019). However, as ADAPT only introduces a minimal number of learnable parameters, these attacking algorithms cannot extract sufficient information from gradients to reconstruct participants’ local data, which gives our ADAPT a significant privacy advantage over traditional federated learning algorithms. Figure 3 presents examples of DLG for our model, where there does not appear any meaningful information corresponding to the input.
Another potential way to expose privacy is decoding the trained text prompts, which might contain some statistical information of participants. However, our experiments showcase that this is difficult as well. we follow CoOp (Zhou et al. 2021) to decode each text prompt by finding a standard vocabulary word with minimum Euclidean distance to it in the embedding space, and summarize the interpretation results for DomainNet in Table 1. It shows that our prompts tend to capture some high-level and abstract semantics that are difficult to be summarized to standard natural words.
Figure 3: Examples of reconstructed images produced by Deep Leakage from Gradient (Zhu, Liu, and Han 2019). It shows that such gradient attack algorithms cannot reconstruction meaningful information from our model.
Table 1: Nearest Words of textual prompts learned by ADAPT in DomainNet dataset. N/A means non-Latin characters. It shows that our prompts tend to capture high-level and abstract semantics that are difficult to summarize using standard natural language words found in the dictionary.
# clipart info paint quick real sketch
1 ~ fe N/A N/A ° kd
2 N/A # dng , with
3 lh bh some ? N/A N/A
4 and N/A lh N/A the pjf
Experiments